what is the Price of U.S Coupon Bond issued in May 2018 and the Maturity date is May 2020 , with Par value $30 , yearly Coupon payments $5 ,and Interest Rate 7% ?

1
Expert's answer
2021-12-29T12:30:42-0500

Solution:

The price of U.S Coupon Bond = "C\times \frac{1-(1+r)^{-n} }{r} +\frac{F}{(1+r)^{n}}"


Where: C = Coupon = 7"\%" "\times" 30 = 2.1

F = Face or par value = 30

r = Coupon rate = 7%

n = period to maturity = 12yrs


Bond price = "2.1\times \frac{1-(1+0.07)^{-12} }{0.07} +\frac{30}{(1+0.07)^{12}}" = 1 – 7.94 + 13.32 = 6.38


The price of U.S Coupon Bond = 6.38